a. LCompute the product AB for

A = 0 4 0 2 3 1 3 0 1 and B = 1 0 3 1 1 5 2 3 -1

b.) Use your answer in a.) to evaluate det(AB) and compare it to det(A) det(B)

c.) Determine whether or not if det(A+B) is related to det(A) + det(B).


1
Expert's answer
2021-06-01T17:32:46-0400

(a)


"AB=\begin{pmatrix}
 0 & 4 & 0 \\
 2 & 3 & 1 \\
3 & 0 &1
\end{pmatrix}\cdot\begin{pmatrix}
1 & 0 & 3 \\
1 & 1 & 5 \\
2 & 3 & -1
\end{pmatrix}"


"=\begin{pmatrix}
4 & 4 & 20 \\
7 & 6 & 20 \\
5 & 3 &8
\end{pmatrix}"

(b)


"\det(AB)=\begin{vmatrix}
4 & 4 & 20 \\
7 & 6 & 20 \\
5 & 3 &8
\end{vmatrix}"

"=4\begin{vmatrix}
 6 & 20 \\
 3 & 8
\end{vmatrix}-4\begin{vmatrix}
 7 & 20 \\
 5 & 8
\end{vmatrix}+20\begin{vmatrix}
 7 & 6 \\
 5 & 3
\end{vmatrix}"


"=4(48-60)-4(56-100)+20(21-30)=-52"



"\det(A)=\begin{vmatrix}
 0 & 4 & 0 \\
 2 & 3 & 1 \\
3 & 0 &1
\end{vmatrix}"

"=0\begin{vmatrix}
 3 & 1 \\
 0 & 1
\end{vmatrix}-4\begin{vmatrix}
 2 & 1 \\
 3 & 1
\end{vmatrix}+0\begin{vmatrix}
 2 & 3 \\
 3 & 0
\end{vmatrix}"


"=0-4(2-3)+0=4"





"\det(B)=\begin{vmatrix}
1 & 0 & 3 \\
1 & 1 & 5 \\
2 & 3 & -1
\end{vmatrix}"

"=1\begin{vmatrix}
 1 & 5 \\
 3 & -1
\end{vmatrix}-0\begin{vmatrix}
 1 & 5 \\
 2 & -1
\end{vmatrix}+3\begin{vmatrix}
 1 & 1 \\
 2 & 3
\end{vmatrix}"


"=1(-1-15)-0+3(3-2)=-13"


"\det(AB)=-52=\det(A)\det(B)"


(c)


"A+B=\begin{pmatrix}
 0 & 4 & 0 \\
 2 & 3 & 1 \\
3 & 0 &1
\end{pmatrix}+\begin{pmatrix}
1 & 0 & 3 \\
1 & 1 & 5 \\
2 & 3 & -1
\end{pmatrix}"

"=\begin{pmatrix}
1 & 4 & 3 \\
3 & 4 & 6 \\
5 & 3 & 0
\end{pmatrix}"



"\det(A+B)=\begin{vmatrix}
 1 & 4 & 3 \\
 3 & 4 & 6 \\
5 & 3 & 0
\end{vmatrix}"

"=1\begin{vmatrix}
 4 & 6 \\
 3 & 0
\end{vmatrix}-4\begin{vmatrix}
 3 & 6 \\
 5 & 0
\end{vmatrix}+3\begin{vmatrix}
 3 & 4 \\
 5 & 3
\end{vmatrix}"



"=-18-4(-30)+3(9-20)=69"


"\det(A+B)\not=\det(A)+\det(B)"